OpenStudy (anonymous):

can someone help me with this conversion, Due to continental drift, the North American and European continents are drifting apart at an average speed of about 3 cm per year. At this speed, how long will it take for them to drift apart by another 1169 m?

OpenStudy (anonymous):

\[displacement = velocity \times time\]\[\text{distance (metres)} = \text{distance (centimetres)} \times \frac{\text{# metres}}{\text{1 centimetre}}\]How many metres are there in 1 centimetre? There are 0.01